# Entity-aware Cross-lingual Claim Detection for Automated Fact-checking

### v1

@@ Abstract
Identifying claims requiring verification is a critical task in automated fact-checking, especially given the proliferation of misinformation on social media platforms. Despite significant progress in the task, there remain open challenges such as dealing with multilingual and multimodal data prevalent in online discourse. Addressing the multilingual challenge, recent efforts have focused on fine-tuning pre-trained multilingual language models. While these models can handle multiple languages, their ability to effectively transfer cross-lingual knowledge for detecting claims spreading on social media remains under-explored. In this paper, we introduce EX-Claim, an entity-aware cross-lingual claim detection model that generalizes well to handle claims written in any language. The model leverages entity information derived from named entity recognition and entity linking techniques to improve the language-level performance of both seen and unseen languages during training. Extensive experiments conducted on three datasets from different social media platforms demonstrate that our proposed model significantly outperforms the baselines, across 27 languages, and achieves the highest rate of knowledge transfer, even with limited training data.
Introduction
@@
Automated fact-checking is an emerging research task dedicated to combating misinformation, particularly prevalent on social media, and comprises several key stages: claim detection, claim prioritization, evidence retrieval, and claim validation (Zeng, Abumansour, and Zubiaga 2021). While research in the task is progressing rapidly, identifying and validating claims related to global concerns requires a fact-checking pipeline capable of processing claims written in multiple languages. However, developing multilingual solutions for fact-checking research is not straightforward due to the availability of limited training data, especially for low-resource languages (Panchendrarajan and Zubiaga 2024). This necessitates advancements in multilingual fact-checking research to leverage limited training data for developing generalized solutions that can effectively transfer knowledge between languages.
@@
This research focuses on the first component of the fact-checking pipeline: verifiable claim detection. A verifiable claim is defined as a statement expressing facts, excluding personal experience and private knowledge (Panchendrarajan and Zubiaga 2024). Existing works on verifiable claims detection are predominantly focused on monolingual solutions (Prabhakar, Mohtaj, and Möller 2020; Suri and Dudeja 2022; Henia et al. 2021; Hussein, Ghneim, and Joukhadar 2021), with limited attention given to addressing the multilingual nature of the problem. Especially, fine-tuning multilingual language models such as mBERT (Alam et al. 2021b; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Panda and Levitan 2021) and XLM-R (Alam et al. 2021b; Hüsünbeyi, Deck, and Scheffler 2022; Savchev 2022; Alam et al. 2021a) is observed as a common solution. While these models can handle multiple languages, none of them adequately validate the cross-lingual knowledge-transferring capability of these models, particularly in identifying claims written in languages that were not seen during training. To address this challenge, we propose EX-Claim, a cross-lingual claim detection model that can effectively identify verifiable claims regardless of their language. Our model broadens the ability to perform across a diverse set of languages beyond those seen in training by incorporating entity-centric information for knowledge transfer. Our approach leverages the limited multilingual data available in claim detection research to develop a robust cross-lingual model that generalizes beyond the languages seen in the training phase.
@@
  • S1: X visited Denmark to attend ICWSM
@@
Our proposed EX-Claim builds on this intuition by leveraging entity information to effectively identify verifiable claims in a cross-lingual environment. Our model incorporates entity information such as type and popularity derived from named entity recognition (NER) and entity linking (EL). Extensive experiments conducted using three datasets from different social media platforms comprising 27 languages show that EX-Claim significantly outperforms the baselines and reaches the highest knowledge transfer rate between languages seen and unseen in training. We make the following key contributions:
@@
Related Work
A common approach to identifying multilingual claims shared on social media involves fine-tuning pre-trained multilingual language models such as mBERT (Alam et al. 2021b; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Panda and Levitan 2021; Zengin, Kartal, and Kutlu 2021; Hasanain and Elsayed 2022) and XLM-R (Alam et al. 2021b, a). Most of these works (Alam et al. 2021b, a; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Baris-Schlicht, Magnossao de Paula, and Rosso 2021; Du, Gollapalli, and Ng 2022) fine-tune the language model using combined training data of multiple languages and evaluate the capability of the model in identifying claims written in the same set of languages. Very few studies (Panda and Levitan 2021; Zengin, Kartal, and Kutlu 2021; Hasanain and Elsayed 2022) explore the model’s ability to transfer knowledge across language pairs, and these studies are limited to a narrow selection of languages.
Rather than focusing on developing cross-lingual methods, several existing studies have aimed at creating language-specific models to achieve optimal performance. For instance, Hüsünbeyi, Deck, and Scheffler (2022); Savchev (2022); Eyuboglu et al. (2023) fine-tuned models such as mBERT or XLM-R using language-specific training data. With the recent attention to large language models (LLM), Agrestia, Hashemianb, and Carmanc (2022) utilized GPT-3 to fine-tune the model in each language to develop a language-specific claim detection model. Interestingly, the results revealed that transformer-based architectures like BERT remain competitive with, and in some cases outperform the large language models (LLMs). Following these findings, further research (Sawiński et al. 2023; Lewoniewski et al. 2024; Li, Panchendrarajan, and Zubiaga 2024) explored leveraging LLMs for claim detection by modeling the task as a text generation problem. While these models demonstrate superior performance in the English language, their ability to handle multilingual claims for the claim detection task has yet to be explored.
Apart from the direct utilization of multilingual language models, various other techniques have been employed to address the challenges posed by limited training data and the noisy nature of social media data. This includes the extraction of platform-specific textual features (Alam et al. 2021b), data augmentation techniques such as machine translation (Zengin, Kartal, and Kutlu 2021; Savchev 2022) and sampling (Zengin, Kartal, and Kutlu 2021) and multitasking approaches (Baris-Schlicht, Magnossao de Paula, and Rosso 2021; Du, Gollapalli, and Ng 2022). While these techniques are explored in multilingual environments as well, none of them have been shown to improve cross-lingual knowledge-transferring capabilities, which our study performs comprehensively for the first time looking at 27 different languages through the exploration of entity-based knowledge transfer.
@@
Methodology
@@
Cross-lingual Word Embedding
@@
We exploit various multilingual pre-trained language models to extract embeddings from their final layer. Especially, we experiment with XLM-R (Kalyan, Rajasekharan, and Sangeetha 2021), and mBERT (Devlin et al. 2019) to obtain the cross-lingual vector representation of words. XLM-R was trained on CommonCrawl data supporting 100 languages, and mBERT was trained on Wikipedia data containing 104 languages. Both models generate embeddings of size 768 for each tokenized word.
Named Entity Recognition (NER)
@@
Identifying entities mentioned in the text is essential to extract further information such as its type and popularity. We use state-of-the-art NER models WikiNEuRal (Tedeschi et al. 2021) and MultiNERD (Tedeschi and Navigli 2022) for identifying coarse-grained and fine-grained entities respectively. We use the WikiNEuRal model and fine-tune the MultiNERD model with recommended parameters using their source code and annotated dataset. WikiNEuRal recognizes 4 types of coarse-grained named entities and MultiNERD recognizes 15 types of fine-grained named entities. Refer to Appendix A and A for the named entity types and statistics of named entities recognized.
Entity Linking (EL)
@@
We consider an entity popular if it can be linked to a Wikipedia article. Hence, we run EL for all entities identified in the previous stage to determine if an entity has an associated Wikipedia article that can be linked to. To do this, we make use of a state-of-the-art multilingual EL model, mGENERE (De Cao et al. 2022) for linking entities (either coarse-grained or fine-grained). Given an entity-tagged sentence (the start and end of an entity are indicated explicitly with the special tags [START] and [END]), mGENRE outputs a ranked list of $n$ Wikipedia page titles along with their associated log probabilities, from which we make use of the top item with the highest probability to determine if indeed the entity has a relevant Wikipedia article that can be linked to. If the log probability is above a threshold, we consider that entity as popular.
@@
During parameter tuning, we observed that the variation in threshold did not significantly affect the performance of the claim detection models that utilize EL. This could be due to the majority of named entities being recognized as Wikipedia entities, rendering the information redundant for the model. A possible cause of this nature is NER models using Wikipedia as one of the sources to generate the NER training data. Consequently, we set the EL threshold to -0.15 (probability score of 0.861), as this value resulted in a significant drop in the number of entities linked (refer to Figure 5 in Appendix A), ensuring that only highly probable entity links are used to determine the popularity. Refer to Appendix A for the statistics of named entities linked.
Entity Type Embedding
@@
Similar to the word embedding, an entity type can be represented in a dense vector representation, enabling the model to capture entity type information required for the task as a dense vector. This is achieved by adding an embedding layer to the claim detection model which converts each entity type to an entity type vector. Based on the entity type leveraged, we experiment with the following four variants of our EX-Claim model:
  • EX-Claim-C: Uses four coarse-grained named entities recognized and other indicating words that are not named entities.
@@
  • EX-Claim-CP: Uses both coarse-grained NER and the popularity of an entity. Each named entity can be either popular or unpopular, doubling the size of the entity types. This results in 9 entity types (2*coarse-grained NER types + 1).
  • EX-Claim-F: Uses 15 fine-grained named entities recognized and other indicating words that are not named entities.
  • EX-Claim-FP: Uses both fine-grained NER and the popularity of an entity, resulting in 31 entity types (2*fine-grained NER types + 1).
Category Entity Type Criteria
No of.
Entity Types
X-Claim None 0
EX-Claim-C Coarse-grained NER 5
EX-Claim-CP Coarse-grained NER & Entity Linking 9
EX-Claim-F Fine-grained NER 15
EX-Claim-FP Fine-grained NER & Entity Linking 31
Table 1: Entity Type Categories Experimented
@@
Table 1 summarizes the entity type categories experimented in this study. As an ablated variant, we experiment with X-Claim category which does not use any entity-type information to show the effectiveness of utilizing entity-type information for verifiable claim detection. Except for the X-Claim category, each word in the input sequence is identified with an entity type index $et_{i}$ according to one of the entity type categories. This information is fed into the entity type embedding layer to obtain a dense vector representation $ee_{i}$ for each word in the input sequence.
Entity-aware Claim Detection
@@
To attend to the important words indicating the verifiability of claims, we apply scaled dot-product attention (Vaswani et al. 2017) to the projected input sequence $H=(h_{1},h_{2},...h_{n})$ as follows,
@@
Given the true label distribution $\hat{y}\in\mathbb{R}^{C}$, the entity type embedding layer, bidirectional GRU, and attention layer are trained by optimizing the cross-entropy loss as follows,
$L_{CE}=-\sum_{j=1}^{C}\hat{y}_{j}log(y_{j})$ (7)
Experiment Setting
Dataset
Detecting claims across languages, and evaluating the model’s capability in identifying the same claim across multiple languages requires a parallel and multilingual ground truth dataset. We use CheckThat! 2022 (Nakov et al. 2022) dataset for training and testing the claim detection models. This dataset contains tweets related to COVID-19, written in Arabic, Bulgarian, Dutch, English and Turkish. Further, it contains four partitions: Train, Dev, Dev-test, and Test. We used the Train partition of all languages except Arabic for training and used the Dev partition as validation data. The Arabic data was excluded during the training as the initial experiments yielded poor performance in all languages including Arabic. Further, the Dev-Test partition was used to find the optimal parameters.
To further evaluate the capability of the models in identifying claims written in a broad range of unseen languages, we generated synthetic data from CheckThat! Test partition. The synthetic data was generated for 18 unseen global languages. For each unseen language, we randomly sampled 250 data instances from the Test partition of CheckThat! data. Sampled data instances were translated to the target language using Microsoft Azure AI translator (Junczys-Dowmunt 2019). Every time, a new set of 250 data instances was sampled to ensure that the same set of tweets did not appear across the languages.
To evaluate the generalization capability of the proposed model in identifying claims discussing unseen topics as well as unseen language, we report the performance in Kazemi 2021 (Kazemi et al. 2021). This dataset contains WhatsApp tipline and social group messages written in in high-resource languages such as English and Hindi, and low-resource languages such as Bengali, Malayalam, and Tamil. Refer to Appendix A for the statistics of the datasets.
The claim detection models were evaluated using these three datasets gathered from different social media platforms, enabling assessment of generalization of the proposed solution across the following scenarios:
  • Known context and seen languages - Models were tested on the CheckThat! 2022 test partition. Except for Arabic, the rest of the languages were seen by the model during the training.
  • Known context and unseen languages - Models were tested on synthetic data composed of 18 languages that were not present in the training data.
  • Unknown context and seen/unseen languages - Models were tested on the Kazemi 2021 dataset. Except for English, the other four languages are unseen during the training.
Hyperparameters
Hyperparameter Value
Projection Size 256
Learning Rate 3e-5
Entity Embedding Size 128 (C,F), 256 (CP,FP)
Entity Linking Threshold -0.15
Table 3: Hyperparameters Used
Table 3 lists the hyperparameters used. Following (Devlin et al. 2019), we set the maximum length of the input sequence to 128, and use Adam as the optimizer. We train the models for 30 epochs and choose the model with the lowest validation loss as the best model. Initial experiments revealed, that the proposed solution performed better when XLM-R was used as the cross-lingual embedding layer instead of mBERT. Therefore, the rest of the experiments use XLM-R as the cross-lingual word embedding layer in EX-Claim.
@@
Entity Types Categories X-Claim 0.747 $\pm$ 0.005 0.751 $\pm$ 0.005 0.747 $\pm$ 0.005 0.745 $\pm$ 0.005
EX-Claim-C 0.752 $\pm$ 0.005 0.755 $\pm$ 0.005 0.752 $\pm$ 0.005 0.75 $\pm$ 0.005
EX-Claim-CP 0.753 $\pm$ 0.007 0.756 $\pm$ 0.007 0.753 $\pm$ 0.007 0.751 $\pm$ 0.007
EX-Claim-F 0.754 $\pm$ 0.005 0.758 $\pm$ 0.006 0.754 $\pm$ 0.005 0.752 $\pm$ 0.006
EX-Claim-FP 0.755 $\pm$ 0.007 0.759 $\pm$ 0.007 0.755 $\pm$ 0.007 0.754 $\pm$ 0.007
@@
Entity Types Categories X-Claim 0.719 $\pm$ 0.007 0.724 $\pm$ 0.006 0.719 $\pm$ 0.007 0.718 $\pm$ 0.007
EX-Claim-C 0.729 $\pm$ 0.007 0.734 $\pm$ 0.007 0.729 $\pm$ 0.007 0.728 $\pm$ 0.007
EX-Claim-CP 0.727 $\pm$ 0.01 0.733 $\pm$ 0.008 0.727 $\pm$ 0.01 0.727 $\pm$ 0.01
EX-Claim-F 0.73 $\pm$ 0.008 0.735 $\pm$ 0.007 0.73 $\pm$ 0.008 0.729 $\pm$ 0.008
EX-Claim-FP 0.731 $\pm$ 0.009 0.737 $\pm$ 0.008 0.731 $\pm$ 0.009 0.73 $\pm$ 0.009
@@
Entity Types Categories X-Claim 0.788 $\pm$ 0.007 0.817 $\pm$ 0.005 0.788 $\pm$ 0.007 0.789 $\pm$ 0.007
EX-Claim-C 0.786 $\pm$ 0.009 0.806 $\pm$ 0.007 0.786 $\pm$ 0.009 0.789 $\pm$ 0.008
EX-Claim-CP 0.785 $\pm$ 0.007 0.803 $\pm$ 0.007 0.785 $\pm$ 0.007 0.788 $\pm$ 0.007
EX-Claim-F 0.793 $\pm$ 0.008 0.813 $\pm$ 0.006 0.793 $\pm$ 0.008 0.796 $\pm$ 0.008
EX-Claim-FP 0.794 $\pm$ 0.008 0.813 $\pm$ 0.006 0.794 $\pm$ 0.008 0.797 $\pm$ 0.007
@@
Entity Types Categories X-Claim 0.736 $\pm$ 0.006 0.746 $\pm$ 0.006 0.736 $\pm$ 0.006 0.735 $\pm$ 0.007
EX-Claim-C 0.743 $\pm$ 0.007 0.751 $\pm$ 0.006 0.743 $\pm$ 0.007 0.743 $\pm$ 0.007
EX-Claim-CP 0.742 $\pm$ 0.009 0.75 $\pm$ 0.008 0.742 $\pm$ 0.009 0.742 $\pm$ 0.009
EX-Claim-F 0.746 $\pm$ 0.007 0.753 $\pm$ 0.007 0.746 $\pm$ 0.007 0.745 $\pm$ 0.008
EX-Claim-FP 0.747 $\pm$ 0.008 0.754 $\pm$ 0.007 0.747 $\pm$ 0.008 0.746 $\pm$ 0.009

### v3

@@ Abstract
Identifying claims requiring verification is a critical task in automated fact-checking, especially given the proliferation of misinformation on social media platforms. Despite notable progress, challenges remain—particularly in handling multilingual data prevalent in online discourse. Recent efforts have focused on fine-tuning pre-trained multilingual language models to address this. While these models can handle multiple languages, their ability to effectively transfer cross-lingual knowledge for detecting claims spreading on social media remains under-explored. In this paper, we introduce EX-Claim, an entity-aware cross-lingual claim detection model that generalizes well to handle multilingual claims. The model leverages entity information derived from named entity recognition and entity linking techniques to improve the language-level performance of both seen and unseen languages during training. Extensive experiments conducted on three datasets from different social media platforms demonstrate that our proposed model stands out as an effective solution, demonstrating consistent performance gains across 27 languages and robust knowledge transfer between languages seen and unseen during training.
Entity-aware Cross-lingual Claim Detection for Automated Fact-checking
Rrubaa Panchendrarajan, Arkaitz Zubiaga School of Electronic Engineering and Computer Science Queen Mary University of London {r.panchendrarajan, a.zubiaga}@qmul.ac.uk
1 Introduction
@@
Automated fact-checking is an emerging research task dedicated to combating misinformation, particularly prevalent on social media, and comprises several key stages: claim detection, claim prioritization, evidence retrieval, and claim validation (Zeng et al., 2021). While research in the task is progressing rapidly, identifying and validating claims related to global concerns requires a fact-checking pipeline capable of processing claims written in multiple languages. However, developing multilingual solutions for fact-checking research is not straightforward due to the availability of limited training data, especially for low-resource languages (Panchendrarajan and Zubiaga, 2024). This necessitates advancements in multilingual fact-checking research to leverage limited multilingual training data for developing generalized solutions that can effectively transfer knowledge between languages.
@@
This research focuses on the first component of the fact-checking pipeline: verifiable claim detection. A verifiable claim is defined as a statement expressing facts, excluding personal experience and private knowledge (Panchendrarajan and Zubiaga, 2024). Existing works on verifiable claims detection are predominantly focused on monolingual solutions (Prabhakar et al., 2020; Suri and Dudeja, 2022; Henia et al., 2021; Hussein et al., 2021), with limited attention given to addressing the multilingual nature of the problem. Especially, fine-tuning multilingual language models such as mBERT (Alam et al., 2021b; Uyangodage et al., 2021; Panda and Levitan, 2021) and XLM-R (Alam et al., 2021b; Hüsünbeyi et al., 2022; Savchev, 2022; Alam et al., 2021a) is observed as a common solution. While these models can handle multiple languages, none of them adequately validate the cross-lingual knowledge-transferring capability of these models, particularly in identifying claims written in languages that were not seen during training. To address this challenge, we propose EX-Claim, a cross-lingual claim detection model that can effectively identify verifiable claims regardless of their language. Our model broadens the ability to perform across a diverse set of languages beyond those seen in training by incorporating entity-centric information for knowledge transfer. Our approach leverages the limited multilingual data available in claim detection research to develop a robust cross-lingual model that generalizes beyond the languages seen in the training phase.
@@
  • S1: X visited China to attend EMNLP
@@
Our proposed EX-Claim builds on this intuition by leveraging entity information to effectively identify verifiable claims in a cross-lingual environment. Our model incorporates entity information such as type and popularity derived from named entity recognition (NER) and entity linking (EL). Extensive experiments conducted using three datasets from different social media platforms comprising 27 languages show that EX-Claim stands out as the effective model with consistent performance gain across multiple languages and robust knowledge transfer between languages seen and unseen in training. We make the following key contributions:
@@
2 Related Work
A common approach to identifying multilingual claims shared on social media involves fine-tuning pre-trained multilingual language models such as mBERT (Alam et al., 2021b; Uyangodage et al., 2021; Panda and Levitan, 2021; Zengin et al., 2021; Hasanain and Elsayed, 2022) and XLM-R (Alam et al., 2021b, a) and developing language-specific models to achieve optimal performance Hüsünbeyi et al. (2022); Savchev (2022); Eyuboglu et al. (2023). Most of these works (Alam et al., 2021b, a; Uyangodage et al., 2021; Baris-Schlicht et al., 2021; Du et al., 2022) fine-tune the language model using combined training data of multiple languages and evaluate the capability of the model in identifying claims written in the same set of languages. Very few studies (Panda and Levitan, 2021; Zengin et al., 2021; Hasanain and Elsayed, 2022) explore the model’s ability to transfer knowledge across language pairs, and these studies are limited to a narrow selection of languages.
@@ 2 Related Work
With the recent attention to large language models (LLM), Agrestia et al. (2022) utilized GPT-3 to fine-tune the model in each language to develop a language-specific claim detection model. Interestingly, the results revealed that transformer-based architectures like BERT remain competitive with, and in some cases outperform the large language models (LLMs). Following these findings, further research (Sawiński et al., 2023; Lewoniewski et al., 2024; Li et al., 2024) explored leveraging LLMs for claim detection by modeling the task as a text generation problem. While these models demonstrate superior performance in the English language, their ability to handle multilingual claims for the claim detection task has yet to be explored.
Beyond directly using multilingual language models, various methods have been applied to tackle limited training data and noisy social media text. This includes the extraction of platform-specific textual features (Alam et al., 2021b), data augmentation techniques such as machine translation (Zengin et al., 2021; Savchev, 2022) and sampling (Zengin et al., 2021) and multitasking Baris-Schlicht et al. (2021); Du et al. (2022). While these techniques are explored in multilingual settings, none of them have been shown to improve cross-lingual knowledge-transferring capabilities, which our study performs comprehensively for the first time looking at 27 different languages through the exploration of entity-based knowledge transfer.
3 Methodology
@@
3.1 Cross-lingual Word Embedding
@@
We use XLM-R Kalyan et al. (2021) to extract the cross-lingual vector representation of the words. XLM-R was trained on CommonCrawl data supporting 100 languages, and generates embeddings of size 768 for each tokenized word.
3.2 Named Entity Recognition (NER)
@@
Identifying entities mentioned in the text is essential to extract further information such as its type and popularity. We use state-of-the-art multilingual NER model MultiNERD Tedeschi and Navigli (2022) for identifying named entities. As the model wasn’t released by the authors, we fine-tuned the MultiNERD model with recommended parameters using their source code and training data. MultiNERD recognizes 15 types of fine-grained named entities. Refer to Appendix A.1 for the named entity types and NER statistics.
3.3 Entity Linking (EL)
@@
We consider an entity popular if it can be linked to a Wikipedia article. Hence, we run EL for all entities identified in the previous stage to determine if an entity has an associated Wikipedia article that can be linked to. To do this, we make use of a state-of-the-art multilingual EL model, mGENERE De Cao et al. (2022) for linking entities. Given an entity-tagged sentence, mGENERE outputs a ranked list of $n$ Wikipedia page titles along with their associated log probabilities, from which we make use of the top item with the highest probability to determine if indeed the entity has a relevant Wikipedia article that can be linked to. If the log probability is above a threshold, we consider that entity as popular.
@@
During parameter tuning, we observed that the variation in threshold did not significantly affect the performance of the claim detection models that utilize EL. This is likely because most named entities were already recognized as Wikipedia entities, making the additional information redundant—possibly due to the NER model’s training data being partially sourced from Wikipedia. Consequently, we set the EL threshold to -0.15 (probability 0.861), which notably reduced the number of linked entities (refer to Figure 4 in Appendix A.2), ensuring that only highly probable entity links are used to determine the popularity. Refer to Appendix A.2 for the EL statistics.
3.4 Entity Type Embedding
@@
Similar to the word embedding, an entity type can be represented in a dense vector representation, enabling the model to capture entity type information required for the task as a dense vector. This is achieved by adding an embedding layer to the claim detection model, which converts each entity type to an entity type vector. Based on the entity type leveraged, we experiment with the following two variants of EX-Claim:
  • EXN-Claim: Uses 15 named entity types recognized by the NER model, along with the other category for non-entities—resulting in a total of 16 entity types.
@@
  • EXP-Claim: Uses both NER and the popularity of an entity. Each named entity can be either popular or unpopular, doubling the size of the entity types. This results in 31 entity types (2*15 + 1).
@@
As an ablated variant, we experiment with X-Claim category which does not use any entity-type information to show the effectiveness of utilizing entity-type information for verifiable claim detection. Except for the X-Claim category, each word in the input sequence is identified with an entity type index $et_{i}$ according to one of the entity type categories. This information is fed into the entity type embedding layer to obtain a dense vector representation $ee_{i}$ for each word in the input sequence.
3.5 Entity-aware Claim Detection
@@
To attend to the important words indicating the verifiability of claims, we apply scaled dot-product attention (Vaswani et al., 2017) to the projected input sequence $H=(h_{1},h_{2},...h_{n})$ as follows,
@@
Given the true label distribution $\hat{y}\in\mathbb{R}^{C}$, the entity type embedding layer, bidirectional GRU, and attention layer are trained by optimizing the cross-entropy loss.
4 Experiment Setting
4.1 Dataset
We use CheckThat! 2022 Nakov et al. (2022) dataset for training and testing the claim detection models. This dataset contains tweets related to COVID-19, written in 5 languages. Further, it contains four partitions: Train, Dev, Dev-test, and Test. We used the Train partition of all languages except Arabic for training and used the Dev partition as validation data. The Arabic data was excluded during the training as the initial experiments yielded poor performance in all languages including Arabic. Further, the Dev-Test partition was used to find the optimal parameters.
To evaluate the models’ ability to identify claims written in a broad range of unseen languages, we generated synthetic data from the CheckThat! Test partition for 18 unseen languages. For each, we randomly sampled 250 instances and translated them using Microsoft Azure AI translator Junczys-Dowmunt (2019), limiting samples to match the test dataset size. Translation quality was assessed on the FLORES-200 benchmark Team et al. (2022), with a CHRF score of 54.35. Further details are in Appendix B.2.
To evaluate the generalization capability of the proposed model in identifying claims discussing unseen topics as well as unseen language, we report the performance in Kazemi 2021 Kazemi et al. (2021). This dataset contains WhatsApp tipline and social group messages written in 5 languages. Refer to Appendix A.3 for the statistics of the datasets.
@@ 4.1 Dataset
X-Claim 0.747 $\pm$ 0.005 0.751 $\pm$ 0.005 0.747 $\pm$ 0.005 0.745 $\pm$ 0.005
EXN-Claim 0.754 $\pm$ 0.005 0.758 $\pm$ 0.006 0.754 $\pm$ 0.005 0.752 $\pm$ 0.006
EXP-Claim 0.755 $\pm$ 0.007 0.759 $\pm$ 0.007 0.755 $\pm$ 0.007 0.754 $\pm$ 0.007
@@
X-Claim 0.719 $\pm$ 0.007 0.724 $\pm$ 0.006 0.719 $\pm$ 0.007 0.718 $\pm$ 0.007
EXN-Claim 0.73 $\pm$ 0.008 0.735 $\pm$ 0.007 0.73 $\pm$ 0.008 0.729 $\pm$ 0.008
EXP-Claim 0.731 $\pm$ 0.009 0.737 $\pm$ 0.008 0.731 $\pm$ 0.009 0.73 $\pm$ 0.009
@@
X-Claim 0.788 $\pm$ 0.007 0.817 $\pm$ 0.005 0.788 $\pm$ 0.007 0.789 $\pm$ 0.007
EXN-Claim 0.793 $\pm$ 0.008 0.813 $\pm$ 0.006 0.793 $\pm$ 0.008 0.796 $\pm$ 0.008
EXP-Claim 0.794 $\pm$ 0.008 0.813 $\pm$ 0.006 0.794 $\pm$ 0.008 0.797 $\pm$ 0.007
@@
X-Claim 0.736 $\pm$ 0.006 0.746 $\pm$ 0.006 0.736 $\pm$ 0.006 0.735 $\pm$ 0.007
EXN-Claim 0.746 $\pm$ 0.007 0.753 $\pm$ 0.007 0.746 $\pm$ 0.007 0.745 $\pm$ 0.008
EXP-Claim 0.747 $\pm$ 0.008 0.754 $\pm$ 0.007 0.747 $\pm$ 0.008 0.746 $\pm$ 0.009